How can determine which of the functions f(x) = x^2 and g(x) = 2^x, increases faster over their common domain?
Answer by sofiyac(983)

you plug in same values for x and see which one gives you bigger values for y's
lets plug in x=4
first one you get y=4^2=16
second one you get y=2^4=16
seems equal, now lets plug in x=5
first one you get y=5^2=25
second one you get y=2^5=32
Looks like second function wins!